OpenStudy (anonymous):

*HELP been trying for hours! Verify that quadrilateral BCDE is a rhombus with vertices B(-5, 1), C(0, -4), D(5, 1) and E (0, 6) by showing that all four sides are equal. A. BD = CE = square root 50 B. BD = CE = 50 C. BC = CD = DE = EB = 50 D. BC = CD = DE = EB = square root 50 Help Please!

OpenStudy (anonymous):

We have | BD | = \[\sqrt{(-5-0)^{2}+(1-(-4))^{2}} = \sqrt{25+25}= \sqrt{50}\] Similar to | CE | = | DE | = | EB | .= \[\sqrt{50}\] => D is correct .

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

yw :)